OpenStudy (anonymous):

What is the ratio of x = 5/8y

OpenStudy (anonymous):

I don't understand the question... \(\large x=\frac{5}{8}y\) or \(\large x=\frac{5}{8y}\) . What do you mean by 'what is the ratio?'

OpenStudy (anonymous):

Ratio of the equation. the first one. X=5/8y. Please Help. Thanks.

OpenStudy (anonymous):

That doesn't give me any new information. Which of the equations I mentioned is what you're looking at? is it 'X equals five-eighths of Y' or 'X = five over 8 Y?' "Ratio of the equation" doesn't make any sense. If you are just wondering what ratio is present in the equation, then 5:8 is a ratio, but I still don't really know what you're talking about here.

OpenStudy (anonymous):

x = five is to eight Y. and i need to find the ratio of x and y

OpenStudy (anonymous):

Ah. Ok, that's still a little confusing in the wording, but I think I get the nature of your question now. You're starting with \(\large x=\frac{5}{8y}\) and you want to know what \(\large \frac{x}{y} = ?\) If that is the case, then you need to divide both sides of the equation by y to get \[\large \frac{x}{y}=\frac{5}{8y^2}\]

OpenStudy (anonymous):

I'm not sure really what your question actually is, but if it is as I've shown it, then that is the answer.

OpenStudy (anonymous):

Thanks. :)